a cyclist moving with constant velocity of 6.0m/s forward passes a car that is just starting. if the car has constant acceleration of 2.0m/s^2, where and when will the car overtake the cyclist (ie. they are both the same distance from the starting point)
(this is a kinematics question grade 12)
1
Expert's answer
2018-04-19T09:49:08-0400
They are both the same distance from the starting point:
d_cyclist=d_car
vt=(at^2)/2
2v=at
The time is
t=2v/a=2 6.0/2.0=6.0 s.
The distance is
d=vt=(6.0)(6.0)=36 m.